OpenStudy (anonymous):

What is the solution of x/6 - 8 = 7?

OpenStudy (anonymous):

Possible answers: -6 6 15 90

OpenStudy (anonymous):

\[\frac{ x }{ 6 } = 7+ 8\] \[x = 15 \times 6\] It should be easy now?

OpenStudy (anonymous):

Thanks, I have it now. c:

OpenStudy (anonymous):

Sure, thing. :-]